OpenStudy (rayford):

y varies inversely as twice x. When x = 4, y = 1. Find y when x = 5.

OpenStudy (anonymous):

as it varies inversely: y = k /2x where k is a constant find the value of k by replacins x and y in the above by 4 and 1 so 1 = k / 2*4 now you can find the value of y can you do this?

OpenStudy (rayford):

so 1=k/8?

OpenStudy (anonymous):

right so whats k?

OpenStudy (anonymous):

multiply both sides of the equation by 8 will give you value of k

OpenStudy (rayford):

so 8?=k

OpenStudy (anonymous):

k = 8 now replace k by 8to give the relation of y to x y = 8 / 2x

OpenStudy (anonymous):

and to find y when x = 5 just replace x in the above relation by 5 and calculate

OpenStudy (anonymous):

these problems have 2 steps 1 find value of from data given to get the formula relating y to x 2 substitute value of 1 variable to get the other

OpenStudy (anonymous):

* 1 find value of k

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

You can do it like this: